# Amanpreet Singh  Joseph Chee Chang  Chloe Anastasiades  Dany Haddad Aakanksha Naik  Amber Tanaka  Angele Zamarron  Cecile Nguyen  Jena D. Hwang Jason Dunkleberger  Matt Latzke  Smita Rao Jaron Lochner  Rob Evans

### v1

@@ 4.1 Retrieval
We experimented with medium sized embedding models based on top performers on the retriever and ranking tasks of the MTEB (Muennighoff et al., 2022) leaderboard on HuggingFace. Table 3 in Appendix D lists our candidate models. The mxbai-embed-large-v1 (Lee et al., 2024) embeddings performed best over our dev set. Figure 4 validates our choice of quantization method and target Matryoshka dimension for these embeddings. We chose ubinary quantization with no Matryoshka truncation, (indicated by a red circle on the plot) since it satisfied our storage constraints without a large drop in performance. We experimented with ensembling SparseEmbed Kong et al. (2023), embedding cosine similarity, BM25, and chose the latter two (weight split of $(0.6,0.4)$ respectively) based on the results (See Appendix E). The BM25 scores are normalized with min-max scaling before computing the ensemble score.
@@ 4.3 Generation
Model (Size) Latency (sec/query) nDCG @10 mRR
@@
As shown in Table 2, our system outperforms several popular LLMs: Llama 3.1 (Dubey et al., 2024), Claude Sonnet 3.7 (Anthropic, 2024) and GPT-4o Hurst et al. (2024). It even outperforms reasoning models such as Sonnet 3.7 Thinking (Anthropic, 2025), o1-mini (OpenAI, 2024b) and o3-mini (Zhang et al., 2025) overall on the Scholar QA-CS benchmark. This setup lacks any retrieval so the models generate the responses completely from parametric memory. The benchmark rewards attribution and supporting evidence as a measure of trust in the system, so these models score lower overall. The reasoning based models perform better than our system on the rubrics score, which suggests that they may be superior backbones for our system. However, due to the additional reasoning tokens, these models are more expensive and also significantly increase latency.
@@
Claude 3.0 S* 57.4 50.6 OS-GPT-4o 49.3 53.5 25.9
Claude 3.5 S 50.4 46.6 PaperQA2 38.7 51.4 25.3
@@ 4.4 Real-world Usage and User Feedback
For insight into the failure modes, we manually examined the 383 instances of neutral/negative free-form feedback. We identify five categories (listed by frequency): Request for more detail, Incorrect or missing reference, Off-topic answer or misunderstood query, General feedback about quality, Language or format issues (example feedback in Appendix F). We hypothesize that follow-up questions may help address insufficient answer detail and cases with a lack of retrieved documents, while improved retrieval may help address incomplete or incorrect references and off-topic responses.
@@ Appendix I Query Type Analysis
Claude 3.0 S* 63.0 55.4 OS-GPT-4o 54.9 59.9
Claude 3.5 S 57.8 51.3 PaperQA2 43.8 54.1
Claude 3.7 S 68.4 60.8 Perplex. Sonar DR 43.9 56.0
+Thinking 68.3 58.7 STORM 59.2 64.7
GPT-4o* 65.1 54.1

### v2

@@ 4.1 Retrieval
We experimented with medium sized embedding models based on top performers on the retriever and ranking tasks of the MTEB (Muennighoff et al., 2022) leaderboard on HuggingFace. Table 4 in Appendix D lists our candidate models. The mxbai-embed-large-v1 (Lee et al., 2024) embeddings performed best over our dev set. Figure 4 validates our choice of quantization method and target Matryoshka dimension for these embeddings. We chose ubinary quantization with no Matryoshka truncation, (indicated by a red circle on the plot) since it satisfied our storage constraints without a large drop in performance. We experimented with ensembling SparseEmbed Kong et al. (2023), embedding cosine similarity, BM25, and chose the latter two (weight split of (0.6,0.4)(0.6,0.4)( 0.6 , 0.4 ) respectively) based on the results (See Appendix E). The BM25 scores are normalized with min-max scaling before computing the ensemble score.
@@ 4.3 Generation
Model (Size)
Latency
(sec/query)
nDCG
@10
mRR
@@
As shown in Table 2, our system outperforms popular LLMs: Llama 3.1 (Dubey et al., 2024), GPT 4.1 and Claude Sonnet 3.7 (Anthropic, 2024). It even outperforms reasoning models such as Sonnet 3.7 Thinking (Anthropic, 2025), o1-mini (OpenAI, 2024b) and o3-mini (Zhang et al., 2025) overall on the Scholar QA-CS benchmark. This setup lacks any retrieval so the models generate the responses completely from parametric memory. The benchmark rewards attribution and supporting evidence as a measure of trust in the system, so these models score lower overall. The reasoning based models perform better than our system on the rubrics score, which suggests that they may be superior backbones for our system. However, due to the additional reasoning tokens, these models are more expensive and also significantly increase latency.
@@
Claude 3.5 S 50.4 46.6 OS-GPT-4o 49.3 53.5 25.9
Claude 3.7 S 61.5 55.9 PaperQA2 38.7 51.4 25.3
+Thinking 62.7 55.7 Perplex. Sonar DR 38.7 52.8 25.2
@@ 4.4 Real-world Usage and User Feedback
For insight into the failure modes, we manually examined the 383 instances of neutral/negative free-form feedback. Table 3 lists the feedback types we identified along with their counts as of May 2025 (example feedback in Appendix G). We hypothesize that follow-up questions may help address insufficient answer detail and cases with a lack of retrieved documents, while improved retrieval may help address incomplete or incorrect references and off-topic responses.
Category Count
Incorrect or Missing References 126
Off-topic or Misunderstood Query 113
Request for More Detail or Specificity 289
General Feedback on Quality 149
Language or Format Issues 78
Table 3: Feedback Categories and Counts
@@ 6 Conclusion
Limitations
Supplementing the user feedback discussed in subsection 4.4, we would like to outline some limitations of our system and evaluation and our plans to mitigate them as part of fuure work:
  1. (i)
    Ai2 Scholar QA uses proprietary and closed-source LLM as the backbone for our production pipeline. As shown in Table 2, open source models lag behind the proprietary models in our evaluation. However, we are actively experimenting with open-sourced LLMs to replace the closed ones partially or completely in the pipeline. The open-sourced models will be specifically trained to do well on long-form scientific question answering and each of the sub-tasks in our multi-step generation. Further, our code is open-sourced and can easily be used with potentially any available LLM api provider supported by litellm.
  2. (ii)
    We evaluate the answers generated by Scholar QA and compare against other systems on ScholarQA-CS dataset in subsection 4.3. Even though the answer rubrics are collected via human annotation, the evaluation is only limited to questions in the Computer Science domain and further relies completely on an LLM as the evaluator. In ongoing work, we are investigating more accurate benchmarks for evaluating long form scientific answers. Our approach uses real queries posed by users to Scholar QA, and human preference labels over answers from multiple systems in not just Computer Science, but Biomedicine and other scientific domains. These labels can serve as not only for evaluation, but also as training signals for models.
Acknowledgments
We would like to thank the anonymous reviewers for helpful comments, suggestions and feedback on the manuscript. We would also like to acknowledge the Ai2 ScholarQA users for providing constructive feedback that helped us improve the system. Finally, we thank David Albright for helping with the demo video, the Ai2 communications team for their help with user outreach, and Ai2 engineers and researchers for their help with user testing before launch.
@@ Appendix I Query Type Analysis
Claude 3.5 S 57.8 51.3 OS-GPT-4o 54.9 59.9
Claude 3.7 S 68.4 60.8 PaperQA2 43.8 54.1
+Thinking 68.3 58.7 Perplex. Sonar DR 43.9 56.0